OpenStudy (anonymous):

"n" is an integer, and n^2 < 39. What is the greatest possible value of "n" minus the least possible value of "n"? The answer is 12... how do I compute?

OpenStudy (ranga):

Solve the inequality. n is an integer. That means it can take values such as 1, 2, 3, -1, -2, -3, etc. What is the largest n whose square is less than 39?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I tried squaring 6 to get 36 and a few lower than that but it does not compute to 12

OpenStudy (ranga):

Yes n = +6 will square to 36 which is less than 39. If you go to the next integer it will be 7 and the square is 49 which exceeds 39. So n = +6 is the greatest possible value. Now go in the negative direction for n. n = -1, n = -2, etc. What is the lowest value of n whose square is still less than 39?

OpenStudy (anonymous):

In order to get 12 I would have to square a negative to get 24? Because 6 squared is 36 minus the least possible n.

OpenStudy (ranga):

No. Here we are essentially solving the inequality: n^2 < 39. n = +6 will be the largest integer to satisfy the inequality. n = -6 will be the smallest inequality to satisfy the inequality. Greatest possible value of "n" minus the least possible value of "n" = 6 - (-6) = 12

OpenStudy (anonymous):

OOOOOOHHHHH, I was squaring the greatest n to subtract from the least n and that is not the question. !!!! THANK YOU for clarification !!!!

OpenStudy (ranga):

sure. no problem. You are welcome.